OpenStudy (anonymous):

Please help!!! So Confused!! Find f’(x) given f(x)= [(x^2-2x+4)^3((3-x)/(4x^2))^-2]^-5

OpenStudy (amistre64):

i can see why ... this is more of a practice in the art of divination ...

OpenStudy (anonymous):

at first i wanted to start this problem doing a basic derivative, but i noticed a fraction and thought about quotient rule.

OpenStudy (anonymous):

do you mean like a chain rule

OpenStudy (anonymous):

do as much algebra as you can to simplify this annoying beast before you start

OpenStudy (anonymous):

unless this is in the section on logarthmic differentiation, in which case take the log and then simplify using properties of the log but don't just start taking the derivative of what you have, it will be a nightmare

OpenStudy (anonymous):

using this formulae differentate it with respect to x, f'(x)=(g(x)h'(x)-h(x)g'(x))/(g(x))^2.....where g(x)=(4x^2))^-2]^-5 and h(x)= [(x^2-2x+4)^3((3-x)).... and here an power funtion is also present.. then differentiate them by using power function.. ANOTHER METHOD: taking both side natural log and differentiate on both side and break this complex function by using log property and differentiate them very easily...

OpenStudy (anonymous):

thanks that is a lot easier to understand.